OpenStudy (anonymous):

a jar contains 20 yellow marbles and 55 green marbles and 25 purple marbles. you pick one from the jar at random what is the theoretical probability of picking purple??? percentages

OpenStudy (anonymous):

Purple marbles = 25 Total marbles = 100 25/100 = 1/4 = 25%